# On the magnetic field evolution of interplanetary coronal mass ejections from 0.07 to 5.4 au

### v1

@@ 2.2 An example ICME observed by Solar Orbiter
Figure 2 presents an example of an ICME identified in our catalog using Solar Orbiter magnetic field (MAG instrument, horbury_2020) and proton data (Solar Wind Plasma Analyser (SWA) instrument, owen2020). Here, the ICME start time is given by the shock arrival time, defined by the first data point of an increase in proton speed, density and total magnetic field. The MO start time marks the beginning of the MO interval, which in this case is clearly visible in the Solar Orbiter example as the magnetic field starts to rotate smoothly and the total field jumps to higher values. The proton density and temperature are lower during the MO interval for this event, which is often the case due to expansion. However, in this example, the expected speed signatures are not so obvious, and the density and temperature changes do not perfectly match the magnetic field changes, in particular at the beginning of the MO.
@@
This event serves to demonstrate that our approach is to focus on the magnetic obstacle as identified in the magnetic field data, because magnetometer data are available for all spacecraft, and the magnetic field data often give clear indications of the boundary times by the discontinuities that are present. The fact that the derived plasma times are not always coincident with the magnetic field signatures add to the subjectivity in setting ICME boundaries. From our experience, for many events, the front boundary of the MO is often not so clear, and different catalogs can differ in setting this time (kay2025collection), and sometimes the MO end time is more confidently identifiable than the MO start time, before a transition to the ambient wind occurs.
2.3 Parker Solar Probe events
@@
A major novel data source we explore in this study are events observed close to the Sun by the NASA PSP spacecraft, of which we consider magnetic field (FIELDS; bale2016_fields) and bulk solar wind plasma (Solar Wind Electrons Alphas and Protons, SWEAP; kasper2016_sweap) data. Up to 2025 April 30, which is the end of the PSP/FIELDS dataset we have available at the time of writing, 12 ICMEs have been observed at distances <0.29<0.29 au, and 6 at <0.23<0.23 au. Such close distances to the Sun have never before been reached by any other mission.
@@
Figure 3 shows the magnetic field observations measured by the FIELDS instrument for these 6 events. The closest CME event observed in situ so far is observed on 2022 September 5 (Figure 3a), subject of many studies (e.g. romeo2023_sep2022; paouris2023_sep2022; davies2024; trotta2024; riley2025_sep2022). The CME observed on 2022 June 2 (Figure 3b) was studied by braga2024_june2022, and the shock of an event on 2023 March 13 was analyzed in jebaraj2024. In addition to these known events already discussed in the literature, we have added another three novel events, which we briefly discuss below.
@@
Figure 3c shows the magnetic field data from an event on 2021 April 30, which was observed at 0.0892 au and consists of a shock, with a jump in speed from <200<200 km s-1 to >300>300 km s-1 (not shown), and a magnetic obstacle directly following the shock. It is clearly connected to an eastward directed CME appearing at 2021 April 30 00 UT in SOHO/LASCO/C3. This CME impact happened right after the solar flyby in which PSP first entered below the critical Alfvén surface on 2021 April 28 (kasper2021alfven).
@@
The ICME on 2024 December 22 (Figure 3d) is a weaker event with respect to the general rise in total magnetic field strength, which nonetheless shows a sharp discontinuity in the total field followed by a longer elevation throughout the magnetic obstacle, during which the bipolar $B_{N}$ component changes sign from north to south and the $B_{T}$ component is unipolar, consistent with a low inclination flux rope (bothmer1998).
@@
Finally, the event on 2023 March 13 (Figure 3f) is one of the clearest ICME examples of this subset, along with the 2022 September 5 and 2022 June 2 events. PSP reached a minimum heliocentric distance of 0.2187 au during the ICME observations. It features a very clear shock and strongly elevated total magnetic field in the MO, although it includes a high $B_{R}>0$ component, signaling a more flank-like encounter of PSP with the ICME structure (the shock was studied by jebaraj2024).
@@ 3 Results
3.1 Magnetic field evolution
@@
Previous studies have found different power laws over varying heliocentric distance ranges, in similar formats to Equation 1, with global expansion power law exponents lower within 1 au, and higher beyond 1 au. For example, gulisano2010 considered the mean magnetic field strengths within ICMEs using Helios 1 and 2 observations to obtain an exponent of $k=-1.85\pm 0.07$, covering a heliocentric distance range of 0.3–1 au. This is a much steeper expansion rate than that found by studies using events observed by Ulysses and Juno, covering a heliocentric distance range of 1–5.4 au, with exponents of $k=-1.29\pm 0.12$ (ebert2009bulk), $k=-1.21\pm 0.09$ (richardson2014identification), and $k=1.24\pm 0.43$ (davies2021_juno).
By combining different spacecraft datasets of ICMEs, previous power law exponents over wider heliocentric distance ranges can be determined. For example, catalogs of ICMEs observed by Ulysses extended their power law relationships between 0.3–5.4 au: richardson2014identification included 103 ICMEs observed by the Helios 1 and 2 spacecraft in addition to ICMEs identified at ACE near 1 au in the richardson2010near catalog to find $k=-1.38\pm 0.03$, liu2005statistical found $k=-1.40\pm 0.08$ by including ICMEs identified at Helios, Wind, and ACE, and wang2005characteristics found k=1.52k=-1.52 by including ICMEs identified at Helios, the Pioneer Venus Orbiter and ACE. Similarly, davies2021_juno combined Juno ICME observations with a previous version of the ICMECAT including entries observed by MESSENGER, Venus Express, STEREO-A, STEREO-B, Wind, and MAVEN to determine a power law exponent of $k=-1.66\pm 0.04$ for the flux rope mean magnetic field over the full 0.3–5.4 au heliocentric distance range.
More recently, salman_2024_psp_cat analyzed PSP ICME events between 0.23–0.83 au observed up to mid 2022, before most of the ICME observations of PSP very close to the Sun were made. They find a k=1.21k=-1.21 for the mean magnetic obstacle, which is lower than previous studies for the inner heliosphere (gulisano2010; good2019; Salman_2020). Our aim in this study is now to investigate whether a single power law can accommodate the full distance range from 0.07 to 5.4 au covered in the ICMECAT and whether a connection to solar magnetic field strengths is possible.
@@
[…]
where $R$ is the heliocentric distance, and $B_{0}$ is the power law constant, in our case the magnetic field strength at $R=1$ au. We perform the fitting using the python function scipy.optimize.curve fit. Several algorithms were tested and yielded consistent results. The final results were calculated using the Levenberg-Marquardt algorithm, which provides estimates of the $3\sigma$ spread in the fit parameters. Figure 4a shows the mean magnetic field in the magnetic obstacle $\langle B_{MO}(R)\rangle$ as a function of heliocentric distance $R$. The resulting power law fit is:
[…]
@@
In order to determine whether the innermost ICME observations by PSP (<0.23<0.23 au) are dominantly influencing the results of the power law fits, we repeat the fitting analysis in log-log space with a linear fit. The result is shown in Figure 4b for all events <1.02<1.02 au. There is only a minor change in the power law constant $B_{0}$, but the power law exponent stays almost the same at k=1.58k=-1.58. This means that for the inner heliosphere (<1<1 au) this power law exponent is stable, regardless of the method. Extending the log-log fit to the entire heliocentric distance, as shown in Figure 4c, results in a slightly different exponent of k=1.47k=-1.47, which is consistent with the influence of a weaker $k$ for the outer heliosphere (e.g. gulisano2010; davies2021_juno).
3.2 Relation to solar magnetic field strengths
@@
With a stable power law being able to describe the average CME magnetic field evolution over a wide heliocentric distance range, from 0.07 to 5.42 au, we now examine how the field strengths connect to solar observations, given that the smallest distance of PSP during an observation of a magnetic obstacle was only 0.0685 au (or 14.72 $R_{\odot}$) for the 2022 September 5 CME. For context, this distance lies approximately within the same range as the Alfvén surface as determined from previous PSP solar approaches (16–20 $R_{\odot}$, kasper2021alfven).
Figure 5 shows the total magnetic field magnitude observed by PSP/FIELDS (October 2018 to 30 April 2025) and Solar Orbiter/MAG (April 2020 to 31 May 2025) throughout the mission durations so far, and we indicate the mean magnetic field of the ICME MOs observed within this distance range as colored dots corresponding to the observing spacecraft. Extending the power law for the mean MO magnetic fields to $1~R_{\odot}$ gives a value of about $5\times 10^{4}$ nT, or 0.50 Gauss (1 Gauss $10^{5}$ nT). From studies of the quiet Sun, these average values are known to be around 46 Gauss, or $4.6\times 10^{6}$ nT (e.g. trellesarjona021), i.e. 2 magnitudes higher than implied by the ICME MO power law. For active regions, from which many CMEs originate, the discrepancy is even larger: The average magnetic field strengths of sunspots are on the order of kiloGauss, with averages of 2000 Gauss or $2\times 10^{8}$ nT (e.g. livingston2006_sunspot). This demonstrates that a single power law cannot simultaneously describe both near-Sun and interplanetary magnetic field strengths relevant for ICMEs.
@@
[…]
@@
[…]
@@ 4 Summary and Conclusions
The criteria for an ICME event to be included in the catalog requires a magnetic obstacle (MO) to be present. This choice is motivated by the goal of improving our understanding of the 3D magnetic structure of CME magnetic obstacles. As of today, this is directly only accessible via the 1-D in situ measurements of spacecraft through the enormous CME flux rope structures. MOs often display signatures of flux ropes, featuring smooth rotations of the magnetic field, leading to the strongest geomagnetic storms (zhang2007). While individual boundaries of the MO by different human observers can vary considerably, they are unlikely to influence the results presented here on general CME evolution.
@@
The catalog now includes 12 ICMEs observed by PSP at distances <0.29<0.29 au for the first time. These events were observed by PSP/FIELDS and SWEAP between 2018 and April 2025. With these observations we extend the work of salman_2024_psp_cat to even lower heliocentric distances. Using this expanded dataset from 0.07 to 5.4 au, we show that the magnetic field decay of interplanetary coronal mass ejections and their magnetic obstacles can be described well with a power law with exponent $k$ ranging from k=1.47k=-1.47 to k=1.57k=-1.57, depending on the fitting method and the parameter (i.e. mean or maximum magnetic field strength) the power law fit was applied to.
@@
However, extending this power law to coronal and photospheric distances shows a clear underestimation of these magnetic field strengths of several magnitudes, which we propose to address by introducing multipole power laws with exponents $k_{1}=-1.57$ and $k_{2}=-6$ (for photospheric active region field strengths) and $k_{2}=-4$ (for photospheric quiet Sun field strengths). In comparison, from theory, quadrupole magnetic fields would fall off with distance as k=4k=-4. We point out that a closer look into how the magnetic fields in active regions drop off with distance for individual ICME events that are observed by PSP close to the Sun would be very worthwhile, but a closer inspection is out of scope for this study. While our results provide a statistical description of CME magnetic field evolution, we emphasize that individual events can deviate significantly in their field decay exponents, for instance due to interaction with high speed streams or other ICMEs (e.g. Davies_2021_solo; salman_2024_psp_cat).
@@
Power laws as presented here are useful for a wide range of space weather applications, from Faraday rotation methods (e.g. kooi2022_faraday) to flux rope models (e.g. isavnin_2016; weiss_2021_fit; weiss2024distorted; sarkar2024). They also prove particularly useful for ICME magnetic field modeling of in situ observations of ICMEs that are detected near the Sun–Earth line. If a spacecraft far upstream of Earth detects an ICME (at sub-L1 distances), it can transmit the observed in situ magnetic field data to Earth, which results in longer forecast lead times compared to those possible with spacecraft impacted by the ICME at the L1 point. The application of power laws to the sub-L1 magnetic field data then allows to model the field magnitude of the ICME at the distance of Earth, which is useful for improving CME space weather forecasts. This has recently been demonstrated with STEREO-A and Solar Orbiter (laker2024; weiler2025; davies2025_solo), with first attempts going back to lindsay1999 and kubicka2016. Mission concepts that use spacecraft on distant retrograde orbits (DRO, henon1969; perozzi2017), which allow spacecraft as part of a fleet to temporarily reside near the Sun–Earth line, so far include ESA HENON, planned to be launched in 2026 (cicalo2025henon), and ESA SHIELD (as of 2025 in an early planning phase). They build on the Space Weather Diamond (cyr2000) and MIIST mission concepts (lugaz2024ssr). Further studies need to be undertaken to better understand the CME magnetic field evolution in detail with multi-spacecraft observations from 0.8 to 1.0 au, as these are the distances spacecraft on DROs can be placed on to find an optimal trade-off between forecast lead time, accuracy and number of spacecraft needed for such an orbital configuration.
@@
The ICMECAT living catalog is available online for the research community at https://helioforecast.space/icmecat and the filesharing platform figshare, with the latest version available at https://doi.org/10.6084/m9.figshare.6356420. For this study, the ICMECAT v2.3, updated on 2025 October 15, version 24 on figshare, has been used. The python code for producing the figures and results for this paper is available at https://github.com/cmoestl/icmecat_paper_2026 (release v1.0). This work is supported by ERC grant (HELIO4CAST, 10.3030/101042188). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Solar Orbiter magnetometer operations are funded by the UK Space Agency (grant UKRI943). T. H. is supported by STFC grant ST/W001071/1. The PSP/FIELDS experiment was developed and is operated under NASA contract NNN06AA01C. D.H. was supported by the German Ministerium für Wirtschaft und Klimaschutz and the German Zentrum für Luft‐ und Raumfahrt under contract 50QW2202.

### v2

@@ II.2 An example ICME observed by Solar Orbiter
Figure 2 presents an example of an ICME identified in our catalog using Solar Orbiter magnetic field (MAG instrument, Horbury et al., 2020) and proton data (Solar Wind Plasma Analyser (SWA) instrument, Owen et al., 2020). Here, the ICME start time is given by the shock arrival time, defined by the first data point of an increase in proton speed, density, temperature, and total magnetic field. The MO start time marks the beginning of the MO interval, which in this case is clearly visible in the Solar Orbiter example as the magnetic field starts to rotate smoothly and the total field jumps to higher values. The proton density and temperature are lower during the MO interval for this event, which is often the case due to expansion. However, in this example, the expected speed signatures are not so obvious, and the density and temperature changes do not perfectly match the magnetic field changes, in particular at the beginning of the MO.
@@
This event serves to demonstrate that our approach is to focus on the magnetic obstacle as identified in the magnetic field data, because magnetometer data are available for all spacecraft, and the magnetic field data often give clear indications of the boundary times by the discontinuities that are present. The fact that the derived plasma times are not always coincident with the magnetic field signatures add to the subjectivity in setting ICME boundaries. From our experience, for many events, the front boundary of the MO is often not so clear, and different catalogs can differ in setting this time (Kay et al., 2026). Sometimes the MO end time is more confidently identifiable than the MO start time, before a transition to the ambient wind occurs.
II.3 Parker Solar Probe events
@@
A major novel data source we explore in this study are events observed close to the Sun by the NASA PSP spacecraft, of which we consider magnetic field (FIELDS; Bale et al., 2016) and bulk solar wind plasma (Solar Wind Electrons Alphas and Protons, SWEAP; Kasper et al., 2016) data. Up to 2025 April 30, which is the end of the PSP/FIELDS dataset we have available at the time of writing, 12 ICMEs have been observed at distances <0.29<0.29 au, and 6 at <0.23<0.23 au. Such close distances to the Sun have never before been reached by any other mission.
@@
Figure 3 shows the magnetic field observations measured by the FIELDS instrument for these 6 events. The closest CME event observed in situ so far is observed on 2022 September 5 (Figure 3a), subject of many studies (e.g. Romeo et al., 2023; Paouris et al., 2023; Davies et al., 2024; Trotta et al., 2024; Riley et al., 2025). The CME observed on 2022 June 2 (Figure 3b) was studied by Braga et al. (2024), and the shock of an event on 2023 March 13 was analyzed in Jebaraj et al. (2024). In addition to these known events already discussed in the literature, we have added another three novel events, which we briefly discuss below.
@@
Figure 3c shows the magnetic field data from an event on 2021 April 30, which was observed at 0.0892 au and consists of a shock, with a jump in speed from <200<200 km s-1 to >300>300 km s-1 (not shown), and a magnetic obstacle directly following the shock. It is clearly connected to an eastward directed CME appearing at 2021 April 30 00 UT in SOHO/LASCO/C3. This CME impact happened right after the solar flyby in which PSP first entered below the critical Alfvén surface on 2021 April 28 (Kasper et al., 2021).
@@
The ICME on 2024 December 22 (Figure 3d) is a weaker event with respect to the general rise in total magnetic field strength, which nonetheless shows a sharp discontinuity in the total field followed by a longer elevation throughout the magnetic obstacle, during which the bipolar $B_{N}$ component changes sign from north to south and the $B_{T}$ component is unipolar, consistent with a low inclination flux rope (Bothmer and Schwenn, 1998).
@@
Finally, the event on 2023 March 13 (Figure 3f) is one of the clearest ICME examples of this subset, along with the 2022 September 5 and 2022 June 2 events. PSP reached a minimum heliocentric distance of 0.2187 au during the ICME observations. It features a very clear shock and strongly elevated total magnetic field in the MO, although it includes a high $B_{R}>0$ component, signaling a more flank-like encounter of PSP with the ICME structure (the shock was studied by Jebaraj et al., 2024).
@@ III Results
III.1 Magnetic field evolution
@@
Previous studies have found different power laws over varying heliocentric distance ranges, in similar formats to Equation 1 (see below), with global expansion power law exponents lower within 1 au, and higher beyond 1 au. For example, Gulisano et al. (2010) considered the mean magnetic field strengths within ICMEs using Helios 1 and 2 observations to obtain an exponent of $k=-1.85\pm 0.07$, covering a heliocentric distance range of 0.3–1 au. This is a much steeper expansion rate than that found by studies using events observed by Ulysses and Juno, covering a heliocentric distance range of 1–5.4 au, with exponents of $k=-1.29\pm 0.12$ (Ebert et al., 2009), $k=-1.21\pm 0.09$ (Richardson, 2014), and $k=1.24\pm 0.43$ (Davies et al., 2021b).
By combining different spacecraft datasets of ICMEs, previous power law exponents over wider heliocentric distance ranges can be determined. For example, catalogs of ICMEs observed by Ulysses extended their power law relationships between 0.3–5.4 au: Richardson (2014) included 103 ICMEs observed by the Helios 1 and 2 spacecraft in addition to ICMEs identified at ACE near 1 au in the Richardson and Cane (2010) catalog to find $k=-1.38\pm 0.03$, Liu et al. (2005) found $k=-1.40\pm 0.08$ by including ICMEs identified at Helios, Wind, and ACE, and Wang et al. (2005) found k=1.52k=-1.52 by including ICMEs identified at Helios, the Pioneer Venus Orbiter and ACE. Similarly, Davies et al. (2021b) combined Juno ICME observations with a previous version of the ICMECAT including entries observed by MESSENGER, Venus Express, STEREO-A, STEREO-B, Wind, and MAVEN to determine a power law exponent of $k=-1.66\pm 0.04$ for the flux rope mean magnetic field over the full 0.3–5.4 au heliocentric distance range.
More recently, Salman et al. (2024) analyzed PSP ICME events between 0.23–0.83 au observed up to mid 2022, before most of the ICME observations of PSP very close to the Sun were made. They find a k=1.21k=-1.21 for the mean magnetic obstacle, which is lower than previous studies for the inner heliosphere (Gulisano et al., 2010; Good et al., 2019; Salman et al., 2020). Our aim in this study is now to investigate whether a single power law can accommodate the full distance range from 0.07 to 5.4 au covered in the ICMECAT and whether a connection to solar magnetic field strengths is possible.
@@
$B_{MO}(R)~\text{[nT]}=B_{0}\times R~\text{[au]}^{k},$ (1)
where $R$ is the heliocentric distance, and $B_{0}$ is the power law constant, in our case the magnetic field strength at $R=1$ au. We perform the fitting using the python function scipy.optimize.curve fit. Several algorithms were tested and yielded consistent results. The final results were calculated using the Levenberg-Marquardt algorithm, which provides estimates of the $3\sigma$ spread in the fit parameters using the covariance matrix. These errors should be understood as an uncertainty in the overall power law description, while individual events can differ strongly from the range given within the error bars (e.g. Salman et al., 2020; Davies et al., 2022). Figure 4a shows the mean magnetic field in the magnetic obstacle $\langle B_{MO}(R)\rangle$ as a function of heliocentric distance $R$. The resulting power law fit is:
$\langle B_{MO}(R)\rangle~\text{[nT]}=(10.72\pm 0.58)\times R~\text{[au]}^{-1.57\pm 0.02}.$ (2)
@@
In order to determine whether the innermost ICME observations by PSP (<0.23<0.23 au) are dominantly influencing the results of the power law fits, we repeat the fitting analysis in log-log space with a linear fit. This makes the range of the dataset smaller so the largest values will not influence the fit results as much, given the steep rise to higher magnitudes close to the Sun. The result is shown in Figure 4b for all events <1.02<1.02 au (1524 events). There is only a minor change in the power law constant $B_{0}$, but the power law exponent stays almost the same at k=1.58k=-1.58. This means that for the inner heliosphere (<1<1 au) this power law exponent is stable, regardless of the method. Extending the log-log fit to the entire heliocentric distance, as shown in Figure 4c, results in a slightly different exponent of k=1.47k=-1.47. This means that there is indeed a slight tendency for the PSP observations close to the Sun to dominate the fit results when using a wider distance range, as including events >1>1 au makes $k$ smaller compared to the fit in linear space (k=1.58k=-1.58). This is consistent with a well-known weaker $k$ in the outer heliosphere (e.g. Gulisano et al., 2010; Davies et al., 2021b). Indeed, when we create a fit in log-log space to all events >1>1 au (448 events) we find $k=-1.32\pm 0.04$.
The results for all $k$ parameters for different methods and distance ranges are summarized in Table 3. In summary, for practical applications, our analysis results in a scaling law exponent for the mean MO magnetic field in the inner heliosphere (<1<1 au) as $k=-1.57\pm 0.03$, and as $k=-1.33\pm 0.22$ for the outer heliosphere (>1>1 au).
Table 3: Fit results for the exponent $k$ using different methods and heliocentric distance domains for the mean total magnetic field in the magnetic obstacle.
Distance range in au $k$ linear fit $k$ log-log fit number of events
0.07–1.02 $1.57\pm 0.03$ $1.58\pm 0.03$ 1524
0.07–5.42 $1.57\pm 0.02$ $1.47\pm 0.02$ 1972
1.02–5.42 $1.33\pm 0.22$ $1.32\pm 0.04$ 448
III.2 Relation to solar magnetic field strengths
@@
With a stable power law being able to describe the average CME magnetic field evolution over a wide heliocentric distance range, from 0.07 to 5.42 au, we now examine how the field strengths connect to solar observations, given that the smallest distance of PSP during an observation of a magnetic obstacle was only 0.0685 au (or 14.72 $R_{\odot}$) for the 2022 September 5 CME. For context, this distance lies approximately within the same range as the Alfvén surface as determined from previous PSP solar approaches (16–20 $R_{\odot}$, Kasper et al., 2021).
Figure 5 shows the total magnetic field magnitude observed by PSP/FIELDS (October 2018 to 30 April 2025) and Solar Orbiter/MAG (April 2020 to 31 May 2025) throughout the mission durations so far, and we indicate the mean magnetic field of the ICME MOs observed within this distance range as colored dots corresponding to the observing spacecraft. Extending the power law for the mean MO magnetic fields to $1~R_{\odot}$ gives a value of about $5\times 10^{4}$ nT, or 0.50 Gauss (1 Gauss $10^{5}$ nT). From studies of the quiet Sun, these average values are known to be around 46 Gauss, or $4.6\times 10^{6}$ nT (e.g. Trelles Arjona et al., 2021), i.e. 2 magnitudes higher than implied by the ICME MO power law. For active regions, from which many CMEs originate, the discrepancy is even larger: The average magnetic field strengths of sunspots are on the order of kiloGauss, with averages of 2000 Gauss or $2\times 10^{8}$ nT (e.g. Livingston et al., 2006). This demonstrates that a single power law cannot simultaneously describe both near-Sun and interplanetary magnetic field strengths relevant for ICMEs.
@@
$\langle B_{MO}(R)\rangle~\text{[nT]}=10.55\times R~\text{[au]}^{-1.57}+2.02\times 10^{-6}\times R~\text{[au]}^{-6}.$ (5)
@@
$\langle B_{MO}(R)\rangle~\text{[nT]}=9.92\times R~\text{[au]}^{-1.57}+2.1\times 10^{-3}\times R~\text{[au]}^{-4}.$ (6)
@@ IV Summary and Conclusions
The criteria for an ICME event to be included in the catalog requires a magnetic obstacle (MO) to be present. This choice is motivated by the goal of improving our understanding of the 3D magnetic structure of CME magnetic obstacles. As of today, this is directly only accessible via the 1-D in situ measurements of spacecraft through the enormous CME flux rope structures. MOs often display signatures of flux ropes, featuring smooth rotations of the magnetic field, leading to the strongest geomagnetic storms (Zhang et al., 2007). While individual boundaries of the MO by different human observers can vary considerably, they are unlikely to influence the results presented here on general CME evolution.
@@
The catalog now includes 12 ICMEs observed by PSP at distances <0.29<0.29 au for the first time. These events were observed by PSP/FIELDS and SWEAP between 2018 and April 2025. With these observations we extend the work of Salman et al. (2024) to even lower heliocentric distances. Using this expanded dataset from 0.07 to 5.4 au, we show that the magnetic field decay of interplanetary coronal mass ejections and their magnetic obstacles can be described well with a power law with exponent $k$ ranging from k=1.47k=-1.47 to k=1.57k=-1.57, depending on the fitting method and the parameter (i.e. mean or maximum magnetic field strength) the power law fit was applied to. We emphasize that our catalog contains only ICMEs with clear magnetic obstacles, and therefore, the results are applicable to this subset of ICME structures.
@@
However, extending this power law to coronal and photospheric distances shows a clear underestimation of these magnetic field strengths of several magnitudes, which we propose to address by introducing multipole power laws with exponents $k_{1}=-1.57$ and $k_{2}=-6$ (for photospheric active region field strengths) and $k_{2}=-4$ (for photospheric quiet Sun field strengths). In comparison, from theory, quadrupole magnetic fields would fall off with distance as k=4k=-4. This novel technique should be understood as a first heuristic approach to illustrate the possibility for deriving physically predictive models with such a parameterization in the future. The current direct comparison between photospheric magnetic field strengths and fields in magnetic obstacles is certainly a limitation. To this end, a closer look into how the (modeled) coronal magnetic fields in CME source regions decrease with distance would be very worthwhile, in particular for individual ICME events that are observed by PSP close to the Sun. We leave a closer inspection for future studies.
Our results provide a statistical description of CME magnetic field evolution, but we emphasize that individual events can deviate significantly in their field decay exponents, for instance due to interaction with high speed streams or other ICMEs (e.g. Davies et al., 2021a; Salman et al., 2024). Multi-point imaging and in situ observations of the same CME event (e.g. Davies et al., 2021a; Palmerio et al., 2024, 2025) are able to give a more detailed picture, and while the numbers of observed events are still relatively low, they are continuously growing with the fleet of spacecraft currently operating.
Power laws as presented here are useful for a wide range of space weather applications, from Faraday rotation methods (e.g. Kooi et al., 2022) to flux rope models (e.g. Isavnin, 2016; Weiss et al., 2021b, 2024; Sarkar et al., 2020, 2024). For studying solar energetic particle (SEP) propagation (e.g. Rodríguez-García et al., 2025; Wijsen et al., 2025), results of flux rope models often influence how SEPs are accelerated, transported and distributed in the ambient solar wind. It could be possible that remote sensing methods, such as combining observations and models of type IV radio bursts (e.g. Husidic et al., 2025), may shed light on the magnetic field strengths at distances below our in situ limit of 0.07 au, to improve our statistics close to the Sun.
@@
Our results prove particularly useful for ICME magnetic field modeling of in situ observations of ICMEs that are detected near the Sun–Earth line. If a spacecraft far upstream of Earth detects an ICME (at sub-L1 distances), it can transmit the observed in situ magnetic field data to Earth, which results in longer forecast lead times compared to those possible with spacecraft impacted by the ICME at the L1 point. The application of power laws to the sub-L1 magnetic field data then allows to model the field magnitude of the ICME at the distance of Earth, which is useful for improving CME space weather forecasts. This has recently been demonstrated with STEREO-A and Solar Orbiter (Laker et al., 2024; Weiler et al., 2025; Davies et al., 2026), with first attempts going back to Lindsay et al. (1999) and Kubicka et al. (2016). Mission concepts that use spacecraft on distant retrograde orbits (DRO, Henon, 1969; Perozzi et al., 2017), which allow spacecraft as part of a fleet to temporarily reside near the Sun–Earth line, so far include ESA HENON, planned to be launched in 2026 (Cicalò et al., 2025), and ESA SHIELD (as of 2025 in an early planning phase). They build on the Space Weather Diamond (Cyr et al., 2000) and MIIST mission concepts (Lugaz et al., 2024). Further studies need to be undertaken to better understand the CME magnetic field evolution in detail with multi-spacecraft observations from 0.8 to 1.0 au, as these are the distances spacecraft on DROs can be placed on to find an optimal trade-off between forecast lead time, accuracy and number of spacecraft needed for such an orbital configuration.
Data availability statement
The ICMECAT living catalog (Möstl et al., 2025) is available online for the research community at https://helioforecast.space/icmecat and on the filesharing platform figshare ashttps://doi.org/10.6084/m9.figshare.6356420 (catalog doi: 10.6084/m9.figshare.6356420). For this study, the ICMECAT v2.3, updated on 2025 October 15, version 24 on figshare, has been used. The python code for producing the figures and results for this paper is available onhttps://github.com/cmoestl/icmecat_paper_2026 (catalog github) as release v1.1, and in this figshare repositoryhttps://doi.org/10.6084/m9.figshare.31421567 (catalog doi: 10.6084/m9.figshare.31421567).
@@
This work is supported by ERC grant (HELIO4CAST, 10.3030/101042188). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Solar Orbiter magnetometer operations are funded by the UK Space Agency (grant UKRI943). T.H. is supported by STFC grant ST/W001071/1. The PSP/FIELDS experiment was developed and is operated under NASA contract NNN06AA01C. D.H. was supported by the German Ministerium für Wirtschaft und Klimaschutz and the German Zentrum für Luft‐ und Raumfahrt under contract 50QW2202.